# NeuroSim V1.5: Improved Software Backbone for Benchmarking Compute-in-Memory Accelerators with Device and Circuit-level Non-idealities

### v1

@@ III-A Overview of Design Methodology
Designing ACIM accelerators requires optimization across multiple abstraction layers - from device characteristics to system architectures. NeuroSim V1.5 provides an integrated framework to evaluate these designs through two main components shown in Fig. 1 working jointly:
@@
    A behavioral simulator that maps and quantizes neural networks for ACIM implementation and evaluates inference accuracy while accounting for hardware non-idealities.
@@
    A hardware analyzer that estimates system-level energy, latency, and area through detailed circuit models of array peripherals and interconnects.
In this update, our first contribution is a comprehensive re-design of the behavioral simulator. Second, we update the hardware analyzer to support the emerging nvCap-based CIM, complete with circuit models for the new charge-domain compute mode. In the behavioral simulator, we have fundamentally enhanced how we model compute-in-memory operations for accuracy estimation. A key approach is that we leverage pre-characterized statistical models derived from SPICE simulations or silicon measurements to ensure accuracy, while utilizing GPU-accelerated PyTorch operations to achieve the throughput needed for large-scale inference simulation. This separation allows us to maintain high fidelity while dramatically improving simulation efficiency.
The behavioral simulator supports two primary simulation modes:
Device expert mode: Model memory array physics directly, simulating conductance variations, stuck faults, and temporal drift to isolate the effect of memory technologies on network accuracy.
Circuit expert mode: Model memory arrays with a statistical approach using pre-characterized noise distributions from SPICE simulations or silicon measurements, enabling rapid evaluation of aggregated non-idealities from devices and circuits.
Through integration with TensorRT’s post-training quantization flow, NeuroSim V1.5 automatically maps pre-trained neural networks to the corresponding ACIM architecture.
The following sections detail the quantization methodology, array modeling approach, and specific techniques for incorporating device and circuit-level non-idealities into the simulation framework.
@@ III-B1 Quantization
NeuroSim V1.5 adopts NVIDIA’s TensorRT quantization infrastructure for converting neural networks from floating-point to integer representations. This differs from previous NeuroSim versions which used in-house WAGE quantization [59] - an integer-training method that requires retraining the network. TensorRT instead provides post-training quantization through automatic operator replacement, supporting quantization of pre-trained networks using operators like linear layers, convolutions, and pooling without modification. TensorRT supports configurable bit widths for both inputs and weights, which we use to evaluate different precision settings (e.g., 8b/8b, 6b/6b).
The framework determines optimal scaling factors through either max or histogram calibration during a brief calibration phase. This phase typically requires only 2 batches of calibration data (e.g., several hundred images) to analyze the distribution of values in both input tensors and weight matrices. Max calibration simply uses the maximum observed value across these tensors to set the scaling factor. Histogram calibration creates a histogram of the observed tensor values and sets the scale factor based on the cumulative distribution function (CDF) - specifically, choosing the value at which the CDF reaches a user-specified percentile (typically 99.99%). This helps exclude outliers that would otherwise force a larger dynamic range. The key advantages of adopting TensorRT over WAGE quantization are direct support for arbitrary pre-trained models without retraining and the automated support for multiple operator types through operator replacement.
@@ III-B2 Network to Array Mapping
Physical constraints of ACIM arrays necessitate careful partitioning of neural network operations. For a linear layer with input dimension $N$ and output dimension $M$, the quantized weight matrix must be mapped to multiple arrays based on both the memory array dimensions and the precision requirements. The number of memory cells required per weight depends on the bit precision of the cells ($b_{cell}$). For example, to represent an 8-bit weight using 1-bit cells, we need $N_{cell}=\lceil 8/1\rceil=8$ cells per weight. This bit-slicing requirement effectively multiplies the width dimension of our mapping by $N_{cell}$. As shown for a linear layer in Fig. 3, these weight components are arranged in adjacent columns, with each column corresponding to a different bit significance.
Memory array size further constrains the mapping. With a memory array height of $R$ rows and $C$ columns, the $N\times M$ weight matrix must be partitioned into $\lceil N/R\rceil$ rows of arrays. The total number of arrays needed to map the weights is equal to $\lceil N/R\rceil\cdot\lceil MN_{cell}/C\rceil$. Mapping convolutional layers to memory arrays requires additional unfolding operations, of which multiple methods can be implemented. For a more detailed description, readers are referred to [40].
The input feature map is unfolded and mapped to the arrays according to the selected input scheme. In Fig. 2 we show the bit-serial operation, where the feature map is processed separately for each bit of the quantized input.
Outputs from vertically-aligned arrays are accumulated, as these represent partial sums from the same output neurons. Outputs from horizontally aligned arrays are shifted and added according to their bit significance, then concatenated to form the complete layer output.
III-B3 Support for Transformers
For transformer architectures, NeuroSim V1.5 adopts a hybrid approach that maps compatible operations to CIM arrays while implementing other operations in the digital domain. The framework maps linear layers (including query, key, value projections and MLP layers) to CIM arrays using the same methodology as standard neural networks. These linear layers comprise a significant portion of the total computation in transformer models. However, dynamic operations such as attention score computation, layer normalization, and activation functions must still be computed.
We elect to compute these dynamic operations in floating-point compute units due to CIM hardware constraints. Non-volatile memory-based CIM is optimal for weight-stationary computations that do not require frequent weight updates. This is due to limited write endurance and slow programming speed common in these devices. Therefore, operators like self-attention that require unique input and weight matrices each inference are not well-suited for non-volatile CIM. Digital CIM implementations using SRAM could potentially handle these dynamic operations, however they would still require careful quantization of the query, key, and value matrices to maintain compatibility with integer-only computation [60]. Recent work on integer-only transformers [51, 52] has demonstrated promising approaches for full integer quantization, though further research is needed to evaluate the accuracy-efficiency tradeoffs of the hardware.
The current hardware estimation portion of NeuroSim does not include circuit models for floating-point functional units needed for self-attention computation, layer normalization, or GELU activation functions. Future versions will explore integrating these components, potentially through integer approximations that maintain accuracy while enabling efficient hardware implementation.
Figure 3: NeuroSim V1.5 flowchart. Device expert mode: user provides device data (memory state variation, stuck-at-faults, state drift). Circuit expert mode: user provides statistical data on MAC outputs from SPICE Monte-Carlo simulations or tapeout measurements.
@@ III-C1 Memory Array Operation
Building on TensorRT’s operator replacement strategy, NeuroSim V1.5 substitutes standard neural network layers with specialized CIM operators that model array-level computation. When a PyTorch model is loaded, linear and convolutional layers are automatically replaced with cim.Linear and cim.Conv modules that partition operations according to array constraints and simulate analog computation as described in Fig. 3.
The framework incorporates two separate simulation modes for modeling non-idealities: device expert mode for measured device-level device variation or calibrated aging models, and circuit expert mode using statistical models derived from SPICE or silicon measurements. These modes cannot be used simultaneously as circuit-level noise inherently captures device-level variations.
@@ III-C2 Device and Circuit Non-Idealities
In device expert mode, integer weights are first mapped to conductance or capacitance states based on the target memory technology. The framework then simulates analog multiplication through Ohm’s law (I = GV) for resistive arrays or charge-based computation (Q = CV) for capacitive arrays. Input values are treated as idealized voltages, and outputs are computed by summing the resulting currents or charges along each column.
@@
For example, when using 4-bit MLC ($b_{cell}=4$) and 8-bit weights ($b_{w}=8$), $N_{cell}=2$ so $i$ ranges from 0 to 1 and the weight scales for each iteration are $2^{i\cdot b_{cell}}=$ $12^{0\cdot 4}=$ 1 and $2^{i\cdot b_{cell}}=$ $162^{1\cdot 4}=$ 16. When using bit-serial mode ($P_{DAC}=1$) to represent 8-bit activations ($b_{in}=8$), $N_{in}=8$ so $j$ ranges from 0 to 7 with input scales increasing by powers of 2. Each sum in equation 3 corresponds to a for-loop in the behavioral simulation. Consequently, the nested loop becomes the critical piece of code to optimize. Using PyTorch’s GPU-accelerated tensor operations, we can calculate the output $y$ of every memory array on the CIM chip design in parallel, which greatly reduces the overall runtime of NeuroSim.
@@
Stuck-at-Faults (SAF): Memory cells can become permanently fixed to their minimum or maximum states due to manufacturing defects or programming failures. This is modeled by randomly setting cells to minimum or maximum states according to user-specified probabilities before weight mapping. Unlike temporal variations, this permanent modification is applied once during initialization.
@@
In circuit expert mode, statistical models are derived from Monte Carlo circuit analysis or silicon measurements. For SPICE-based characterization, users should:
  • Create a complete array model including peripheral circuits in a circuit simulator.
  • Run Monte Carlo simulations incorporating device variations (threshold voltage, device sizing), circuit-level effects (thermal noise, parasitic), and ADC quantization.
  • Generate statistical distributions by collecting multiple simulation results for each possible MAC output value.
For silicon-based characterization:
  1. 1.
    Collect repeated measurements across multiple array columns.
  2. 2.
    Compare expected vs. actual outputs over thousands of trials.
  3. 3.
    Build output distributions capturing combined effects of all noise sources.
The resulting models provide mean and standard deviation pairs for each possible MAC output value. The data can be displayed in a ”Confusion Matrix” like those shown in Fig. 3. Users provide these parameters as comma-separated tuples, with one variation value per output level. Users store these in the file ’output_noise.csv’. During behavioral simulation, the critical loop (Eq. 3 can be skipped, as the circuit behavior is modeled by the statistical output distribution. Therefore, we can calculate ideal partial sums with negligible runtime overhead. After ideal partial sums are calculated, the framework samples from the MAC distributions to efficiently inject noise into the outputs. This statistical approach enables rapid evaluation and SPICE-level accuracy through empirically validated noise characteristics (see case studies in Section IV.D).
TABLE I: Design Space Exploration Parameters
Parameter Range / Values
@@
Network / Dataset ResNet-18 / CIFAR-100
Array Dimensions $32\times 32$ $256\times 256$
Input / Weight Precision 6b, 8b
@@
Memory Cell Precision 1b -- 4b
ADC Precision 3b -- 11b
  • * The RRAM parameters are configured based on Intel’s 22nm FinFET platform [6]
III-D Performance Analysis Framework
To enable accurate power and performance estimation, NeuroSim V1.5 implements a trace-based interface between its behavioral simulation and hardware evaluation components. During inference simulation in PyTorch, the framework saves quantized input and weight values as decimal traces in CSV format (shown in Fig. 1. These traces capture the actual data patterns that would flow through the hardware, accounting for all quantization effects and bit slicing operations.
The hardware estimator (written in C++) automatically generates a floorplan for the entire accelerator based on the network architecture and user-specified parameters. It creates detailed transistor-level circuit models for all components (memory arrays, ADCs, interconnects, peripherals, accumulators, etc.), calibrated using SPICE simulations and data from previous tapeouts [7]. These models enable accurate hardware performance estimation at advanced technology nodes. Using traces from the software component, the estimator calculates detailed energy consumption for each array. The bit-serial input scheme is explicitly modeled - each input value is processed one bit at a time, with the framework tracking energy consumption for each operation. Weight values from the traces are converted to corresponding conductance or capacitance states based on parameters specified in the hardware configuration file.
The hardware estimator uses these models along with device parameters specified in the configuration file to accurately evaluate PPA metrics for nvCap-based accelerators while maintaining compatibility with the existing trace-based interface between behavioral simulation and hardware analysis.
III-E Evaluation Setup
Our experimental evaluation consists of three main components: design space exploration of hardware configurations, inference accuracy analysis under various device non-idealities, and runtime performance comparison across open-source platforms. The DNN algorithms and proposed NeuroSim framework are implemented using PyTorch, with all simulations performed on a single NVIDIA RTX A6000 GPU with 48GB GDDR6.
III-E1 Design Space Exploration
@@
In our design space exploration, we use binary inputs (1b DAC) and we evaluate two ADC precision settings for each configuration:
@@
Excluding device and circuit noise, the inference accuracy is solely affected by the input/weight quantization from floating-point to fixed-point representation and the quantization of array outputs depending on the ADC precision. Design parameters were swept using automated scripts to explore the trade-offs in accuracy and performance considering input/weight/ADC quantization and array size.
When reducing ADC precision below the ideal precision, we adopt a clipping approach, where the sensing margins for each analog output state remain the same regardless of ADC precision, and outputs larger than the precision of the ADC are clipped to the maximum ADC output based on its precision. We find that this method results in comparable accuracy to other, non-linear quantization methods while being the most practical to implement in hardware.
It is important to note that higher precision ADCs require smaller sensing margins between levels, which in practice could lead to increased susceptibility to circuit noise and higher error rates. The design-space analysis focuses solely on quantization effects, as detailed circuit-level noise characterization data is not available for the multi-bit memory configurations under study. For a complete understanding of system-level reliability, future work should incorporate both quantization effects and circuit-level noise measurements.
III-E2 Device Noise Case Studies
For analyzing inference accuracy under device non-idealities, we standardize the hardware configuration to isolate the effects of various noise sources and variations. Table II presents the fixed parameters used across all device non-ideality studies.
TABLE II: Device Non-idealities Case Study Parameters
Parameter Value
Technology Node 22nm
Device RRAM
[…]
Networks / Datasets VGG8 / CIFAR-10
ResNet-18 / CIFAR-100
ResNet-50 / ImageNet 1k
Swin-T / ImageNet 1k
Array Dimensions $128\times 128$
Activated rows 128
Input / Weight Precision 8b / 8b
Input Processing bit-serial
Memory Cell Precision 1b - 4b
ADC Precision* full precision
Device Noise Sources** D2D, SAF, state drift
  • * ADC precision is calculated using equations (6) and (7)
@@ III-E2 Device Noise Case Studies
III-E3 Circuit Noise Case Studies
@@
III-E4 Runtime Evaluation
To characterize the improvements in runtime over previous versions of NeuroSim, we record the runtime of the inference accuracy portion under several key conditions. The main parameters that influence the runtime are the size of the network and dataset used the MLC level, precision of the DAC if one is used, and the input and weight precision.
IV Results and Case Studies
IV-A Design Space Exploration
We demonstrate NeuroSim V1.5’s capability for comprehensive design space exploration using a 22nm RRAM [6] platform as our baseline technology. Following the parameters outlined in Table I, we evaluate the fundamental tradeoffs between inference accuracy and hardware efficiency across architectural choices including array size, precision, and data representation. Our analysis focuses on architectural optimization without device/circuit noise effects, as these effects are technology-specific and addressed separately in subsequent noise modeling studies.
TABLE III: Default 22nm RRAM PPA (ResNet18 / CIFAR100)

### v2

@@ III-A Overview of Design Methodology
Designing CIM accelerators requires optimization across multiple abstraction layers—from device characteristics to system architectures. NeuroSim V1.5 provides an integrated framework to evaluate these designs through two main components (Fig. 1) working jointly:
@@
    A behavioral simulator that maps quantized neural networks to the CIM floorplan and evaluates inference accuracy under hardware non-idealities.
@@
    A hardware analyzer that estimates system-level energy, latency, and area through detailed circuit models of arrays, peripherals, and interconnects.
The behavioral simulator supports two primary simulation modes: Device expert mode: Model memory array physics directly (conductance variations, stuck faults, temporal drift) to isolate memory technology effects on network accuracy. Circuit expert mode: Model memory arrays statistically using pre-characterized noise distributions (SPICE/silicon), enabling rapid evaluation of aggregated device/circuit non-idealities.
Through TensorRT’s post-training quantization, NeuroSim V1.5 automatically maps pre-trained neural networks to the corresponding CIM architecture. Subsequent sections detail quantization, array modeling, and incorporating non-idealities.
@@ III-B1 Quantization
NeuroSim V1.5 uses TensorRT’s post-training quantization, unlike previous WAGE training-based methods [58]. TensorRT automatically replaces operators (linear, convolution, pooling) with configurable precision (e.g., 8b/8b, 6b/6b).
Scaling factors use max or histogram calibration (99.99% CDF percentile, 2 batches).
@@ III-B2 Network to Array Mapping
Physical CIM array constraints require careful neural network operation partitioning. For example, for a linear layer (input $N$, output $M$), the quantized weight matrix maps to multiple arrays based on array dimensions and precision. The number of memory cells per weight depends on cell bit precision ($b_{cell}$). E.g., an 8-bit weight using 1-bit cells needs $N_{cell}=\lceil 8/1\rceil=8$ cells. Shown for a linear layer in Fig. 2, these weight components split among adjacent columns, each column representing a different bit significance.
Memory array size (R rows, C columns) further constrains mapping. An $N\times M$ weight matrix partitions into $\lceil N/R\rceil$ array rows. Total arrays for weights: $\lceil N/R\rceil\cdot\lceil MN_{cell}/C\rceil$. To map 3D convolutional layers to 2D matrices we used im2col-like mapping described in [39].
The input feature map unfolds and maps per selected input scheme. Fig. 2 shows bit-serial operation (feature map processed separately per quantized input bit). Outputs from vertically-aligned arrays accumulate (partial sums for same output neurons). Outputs from horizontally-aligned arrays shift and add by bit significance, then concatenate for complete layer output.
Figure 3: NeuroSim V1.5 noise modeling flowchart. Device expert mode: user provides device data (memory state variation, stuck-at-faults, state drift). Circuit expert mode: user provides statistical data on MAC outputs from SPICE Monte-Carlo simulations or tapeout measurements.
@@ III-C1 Memory Array Operation
NeuroSim V1.5 replaces neural network layers with specialized CIM operators. Runtime improvements stem from: (1) optimized GPU-accelerated MVM simulation in ‘cim.Linear‘ and ‘cim.Conv‘ modules, and (2) efficient TensorRT quantization. Linear and convolutional layers auto-replace with custom modules that partition operations and simulate analog computation (Fig. 3).
Two simulation modes exist: device expert (measured variations, aging models) and circuit expert (SPICE/silicon statistics), used separately to avoid double-counting.
@@ III-C2 Device and Circuit Non-Idealities
In device expert mode, integer weights map to conductance/capacitance states. The framework simulates analog multiplication via Ohm’s law ($I=GV$) for resistive arrays or charge-based computation ($Q=CV$) for capacitive arrays. Inputs are idealized voltages; outputs are summed column currents/charges. Noise reflecting device non-idealities is typically injected at this stage, pre-ADC quantization.
@@
For example, when using 4-bit MLC ($b_{cell}=4$) and 8-bit weights ($b_{w}=8$), $N_{cell}=2$ so $i$ ranges from 0 to 1 and the weight scales for each iteration are $2^{i\cdot b_{cell}}=$ $2^{0\cdot 4}=1$ and $2^{i\cdot b_{cell}}=$ $2^{1\cdot 4}=16$. When using bit-serial mode ($P_{DAC}=1$) to represent 8-bit activations ($b_{in}=8$), $N_{in}=8$ so $j$ ranges from 0 to 7 with input scales increasing by powers of 2. Each sum in equation 3 corresponds to a for-loop in the behavioral simulation. Consequently, the nested loop becomes the critical piece of code to optimize. Using PyTorch’s GPU-accelerated tensor operations, we can calculate the output $y$ of every memory array on the CIM chip design in parallel, which greatly reduces the overall runtime of NeuroSim.
@@
Stuck-at-Faults (SAF): Cells permanently fixed at min/max states due to defects. Modeled by randomly setting cells per user-specified probabilities during initialization, unlike temporal variations.
@@
Circuit expert mode uses statistical models from Monte Carlo SPICE (varying device properties, temperature, voltage) or silicon measurements (repeated column measurements). Both produce mean/standard deviation pairs for each MAC output, visualized as confusion matrices (Fig. 3) and stored in ’output_noise.csv’.
This approach skips the critical loop (Eq. 3), calculating ideal partial sums with minimal overhead, then sampling MAC distributions for noise injection. Enables rapid evaluation with empirically validated characteristics (Section IV-B).
While this statistical approach captures the aggregate impact of these effects, it doesn’t model detailed physics per input/weight. This results in a trade-off in physics-accurate simulations for fast large-scale design exploration. For highly detailed analysis of such effects, tools like CrossSim [30] may be more suitable.
Additionally, while the current implementation samples device variations with spatial-independence, users can model spatially-correlated defects by supplying the desired variations for those areas. This would require supplying the variation data explicitly for the memory cells in the area of interest along with slight modification to the error injection. Such a feature would be useful for post-silicon measured defects and may be explicitly supported in the future.
III-D Performance Analysis Framework
NeuroSim V1.5 uses a trace-based interface for PPA estimation. The behavioral simulator saves quantized inputs/weights as CSV traces (Fig. 1) capturing hardware data patterns.
The C++ PPA estimator auto-generates accelerator floorplans and creates transistor-level circuit models (arrays, ADCs, interconnects, peripherals) calibrated with SPICE and tapeout data [7]. Using traces, it calculates per-array energy with explicit bit-serial modeling. Weights convert to conductance/capacitance states per configuration parameters.
III-E Support for Vision Transformers
Vision transformer (ViT) architectures are supported in NeuroSim V1.5 using a hybrid ACIM/DCIM approach described in Fig. 4. Linear layers (Q, K, V projections, MLPs) use ACIM arrays in weight-stationary mode. Attention score computation ($QK^{T}$) and weighted aggregation use DCIM arrays, as these operations require writing intermediate attention scores as operands for subsequent multiplication—incompatible with NVM write latency and endurance constraints.
The hybrid architecture maintains NeuroSim’s existing tiled hierarchy. Entire tiles are dedicated to either ACIM or DCIM, with identical bit-serial input processing. Multi-head attention maps naturally to this structure: each head has independent projection weights, enabling parallel execution across tiles. Different layers map to different tiles, and layer-level pipelining allows all tiles to operate simultaneously on different images, maximizing hardware utilization. Outputs from each tile are written to the global buffer before being sent to the next layer. The global buffer is sized to fit the appropriate activation data from each tile operating in parallel. Softmax and GELU are implemented via 8-bit lookup tables (LUTs), storing input-output relationships for the quantized operators.
The DCIM array architecture follows [3] with circuit modeling from [54]. NeuroSim V1.5 supports both noise-integrated accuracy simulation and PPA evaluation for ViT models.
Figure 4: NeuroSim V1.5 ViT dataflow. Linear projections execute on ACIM tiles; attention score and aggregation use DCIM tiles. Numbered circles indicate the operation sequence within one transformer block. Layer-level pipelining maps different layers to different tiles, enabling all tiles to execute in parallel on different images.
TABLE I: Design Space Exploration Parameters and Device Non-idealities Case Study Parameters
Parameter Range / Values for DSE Values for Noise Case Studies
@@ III-E Support for Vision Transformers
Memory Cell Precision 1b - 4b
Network / Dataset ResNet-18 / CIFAR-100 VGG8 / CIFAR-10 ResNet-18 / CIFAR-100 ResNet-50 / ImageNet 1k Swin-T / ImageNet 1k
Array Dimensions $32\times 32$ - $256\times 256$ $128\times 128$
Rows Active all 8 / 32 / 128
Input / Weight Precision 6b, 8b 8b
ADC Precision** 3b - 11b full precision
  • * RRAM parameters based on Intel’s 22 nm FinFET platform [6]
  • ** ADC full precision is calculated using equations (6) and (7)
III-F Evaluation Methodology
Our experimental evaluation consists of three main components: design space exploration of hardware configurations, inference accuracy analysis under various device non-idealities, and runtime performance comparison across open-source platforms. The DNN algorithms and proposed NeuroSim framework are implemented using PyTorch, with all simulations performed on a single NVIDIA RTX A6000 GPU with 48GB GDDR6 memory.
III-F1 Design Space Exploration
@@
In our design space exploration, we evaluate two ADC precision settings for each configuration:
@@
Excluding device and circuit noise, the inference accuracy is solely affected by the input/weight quantization from floating-point to fixed-point representation and the quantization of array outputs depending on the ADC precision. Design parameters were swept using automated scripts to explore the trade-offs in accuracy and performance considering input/weight/ADC quantization and array size. Results are shown in Fig. 5 and discussed further in section IV.
When reducing ADC precision below the ideal precision, we adopt a clipping approach where the sensing margins for each analog output state remain the same regardless of ADC precision. Any outputs larger than the precision of the ADC are clipped to the maximum ADC output based on its precision. We find that this method results in comparable accuracy to dynamic quantization methods while also being the most practical to implement in hardware.
III-F2 Device Noise Case Studies
For analyzing inference accuracy under device non-idealities, we standardize the hardware configuration to isolate the effects of various noise sources and variations. Table I presents the fixed parameters used across all device non-ideality studies.
@@ III-F2 Device Noise Case Studies
III-F3 Circuit Noise Case Studies
@@
III-F4 Runtime Evaluation
To characterize the improvements in runtime over previous versions of NeuroSim, we record the runtime of the inference accuracy portion under several key conditions. These include the network size, input resolution, the MLC level, precision of the DAC (if one is used), and the input and weight precision.
TABLE II: Default 22 nm RRAM PPA (ResNet-18 / CIFAR100)